# Cross-Resolution SAR Target Detection Using Structural Hierarchy Adaptation and Reliable Adjacency Alignment

### v1

@@ Abstract
In recent years, continuous improvements in SAR resolution have significantly benefited applications such as urban monitoring and target detection. However, the improvement in resolution leads to increased discrepancies in scattering characteristics, posing challenges to the generalization ability of target detection models. While domain adaptation technology is a potential solution, the inevitable discrepancies caused by resolution differences often lead to blind feature adaptation and unreliable semantic propagation, ultimately degrading the domain adaptation performance. To address these challenges, this paper proposes a novel SAR target detection method (termed CR-Net), that incorporates structure priors and evidential learning theory into the detection model, enabling reliable domain adaptation for cross-resolution detection. To be specific, CR-Net integrates Structure-induced Hierarchical Feature Adaptation (SHFA) and Reliable Structural Adjacency Alignment (RSAA). SHFA module is introduced to establish structural correlations between targets and achieve structure-aware feature adaptation, thereby enhancing the interpretability of the feature adaptation process. Afterwards, the RSAA module is proposed to enhance reliable semantic alignment, by leveraging the secure adjacency set to transfer valuable discriminative knowledge from the source domain to the target domain. This further improves the discriminability of the detection model in the target domain. Based on experimental results from different-resolution datasets, the proposed CR-Net significantly enhances cross-resolution adaptation by preserving intra-domain structures and improving discriminability. It achieves state-of-the-art (SOTA) performance in cross-resolution SAR target detection.
@@
Cross-resolution detection, evidential learning, SAR target detection, scattering structure, unsupervised domain adaptation, uncertainty quantification.
@@ I Introduction
Synthetic Aperture Radar (SAR) achieves high-resolution imaging by actively transmitting and receiving electromagnetic waves. This unique capability allows SAR to operate effectively under adverse weather conditions and regardless of daylight, enabling all-weather and all-day imaging. These attributes have made SAR widely applicable in the field of Earth observation [1]. As one of the primary means of Earth observation, SAR target detection has been a hot topic in both research and applications. In recent years, advancements in SAR sensor technology have significantly improved the spatial resolution of SAR images, transitioning from traditional meter-level resolution to sub-meter resolution. This leap forward has unlocked significant potential for more refined Earth observation applications: sub-meter resolution SAR images can capture more target details, offering substantial advantages in fine-grained tasks such as urban monitoring, disaster assessment, and target recognition.
@@
Domain adaptation is an effective approach to address the aforementioned domain discrepancies and improve the generalization performance on the unseen data [2, 3]. The goal of this technique is to transfer the knowledge learned from the labeled source domain to an unlabeled target domain with distribution differences, ensuring satisfactory performance on the target domain data. In recent years, some advanced researches have leveraged domain adaptation techniques in SAR target detection, focusing on improving cross-modality (e.g. optical-to-SAR [4, 5, 6] and SAR-to-SAR [7, 8, 9]) detection performance. For instance, Zhao et al. [10, 11] achieved optical-to-SAR target detection by employing adversarial learning and feature decomposition to obtain domain-invariant features. Pan et al. [4] introduced an unbalanced adversarial alignment approach to address the issue of misaligned domain features, enabling ship detection from optical to SAR images. However, some researches [12, 13] have shown that adversarial learning-based domain adaptation methods negatively impact feature discriminability. To address this, some studies [13, 8, 9] have incorporated pseudo-label learning, which aligns feature distributions while enhancing feature discriminability. Additionally, other studies [7, 14, 15] have leveraged the physical priors of SAR targets to enhance SAR target detection under the guidance of prior knowledge. Wang et al. [15] utilized sub-aperture decomposition [16] and data augmentation to achieve the knowledge transfer from synthetic data to real-world data. Some researchers [7] have introduced key scattering points of targets to enable center-aware feature adaptation, improving SAR target detection across different SAR sensors. Zhang et al.[14] incorporated Fourier transform features to establish local structural relationships to enable target detection from optical to SAR images.
Although the aforementioned methods have made some progress in cross-domain SAR target detection, applying these black-box style models to cross-resolution SAR target detection still faces some critical challenges:
@@
    A scattering structure distance is proposed to quantify the structure similarity between SAR targets. Leveraging the scattering structure similarity of instances, a Structure-induced Hierarchical Feature Adaptation (SHFA) module is developed. By introducing structural constraints, SHFA module facilitates hierarchical feature adaptation between the source and target instances, enhancing the feature transferability and avoiding structural distortions throughout the adaptation process.
@@
    A Reliable Structural Adjacency Alignment (RSAA) module is proposed to capture reliable adjacency relationships by aligning semantic information between reliable neighbors across domains. RSAA module effectively transfers discriminative capabilities from the source domain to the target domain, significantly improving the model’s performance on the target domain while mitigating error accumulation.
@@
    The proposed method combines SHFA and RSAA modules, achieving significant performance improvements in cross-resolution SAR target detection tasks. Experimental results demonstrate that SHFA not only preserves the intra-domain structures but also significantly enhances the reliability and accuracy of cross-resolution detection.
The remainder of this article is organized as follows. Section II reviews related works. Section III explains how evidence learning is incorporated into SAR detection models. Building on the evidence-enhanced SAR detection framework, Section IV details the proposed adaptation method. Section V presents the experimental results and analysis. Finally, Section VI concludes the paper.
@@ II-A Uncertainty Estimation and Evidential Learning
Uncertainty is a key factor for evaluating the predictive reliability of deep models [17, 18, 19]. Bayesian deep learning [20] and subsequent approximate methods [21, 22] provide theoretical and practical approaches for uncertainty modeling [20]. However, they still face challenges related to high computational cost and low efficiency. Recently, evidence learning [17, 23] has gained significant attention due to its capability to explicitly model and inference with evidence, making it a powerful approach for quantifying uncertainty. Compared to traditional Bayesian methods, evidence learning learns a Dirichlet distribution to simultaneously output predictive confidence and uncertainty, avoiding high computational costs and improving flexibility [19]. Currently, evidential learning has been widely applied to high-reliability tasks such as medical diagnosis [24] and anomaly detection [25]. However, its application in cross-resolution SAR target detection remains largely unexplored. This paper applies evidence learning in cross-resolution SAR detection tasks to model uncertainty and improve adaptation reliability across resolutions.
@@ II-B SAR Target Detection
In SAR target detection, mainstream solutions are divided into one-stage detectors (e.g., RetinaNet [26], YOLO [27, 28], SSD [29]) and two-stage detectors (e.g., Faster RCNN [30], Sparse RCNN [31]). There are many notable advancements based on these paradigms in SAR community. Zhou et al.[32] proposed a Sidelobe-aware detection network to enhance small ship feature representation. Ju et al.[33] introduced a Polar Encoding and Decoding strategy to improve dense regression for rotated target detection, while Zhao et al.[34] improved the rotated target detection by enhancing inter-class separability [35]. Yang et al.[36] proposed a unified dynamic SAR detection framework t boost dense prediction performance in complex scenarios. For two-stage detectors, Cui et al.[37] improved multi-scale ship detection accuracy using attention mechanisms, and [38] addressed nearshore ship detection problem using an anisotropic pyramid-based saliency enhancement algorithm. Some studies [39, 40, 16] have improved target detectors from the perspective of environmental clutters. For example, certain efforts [39] suppressed the sea clutter using frequency domain information, while Dai et al.[40] addressed low-frequency bias and enhanced high-frequency features through multi-scale dynamic learning. Additionally, some researchers[41, 16] applied data transformations or sub-aperture decomposition [16] to optimize input data quality, thereby improving target detection performance.
@@ II-C Unsupervised Domain Adaptation
Unsupervised Domain Adaptation (UDA) is an important research direction in transfer learning, aiming to improve the performance of models on the target domain by leveraging labeled data from the source domain and unlabeled data from the target domain [2]. The core challenge lies in the distribution discrepancy (i.e., domain shift) between the source and target domains. Researchers have proposed various methods to reduce this discrepancy, including distribution alignment (e.g., DAN [2], DANN [3], MADA [42]), pseudo-label generation [43, 13], image translation methods [12, 44, 45], and mean-teacher training methods [46, 47]. UDA is widely applied in fields such as computer vision [48, 3, 46, 49], natural language processing [50], medical image analysis [51, 52], remote sensing data processing [5, 9], and autonomous driving [47]. Despite significant progress, UDA still faces challenges such as interpretability [53], domain structure distortion [2], and unreliable knowledge transfer [19].
@@ II-D UDA for SAR Target Detection
In recent years, UDA has become a popular research topic in SAR target detection. Mainstream methods include image translation [12, 54], adversarial adaptation [55, 56], and pseudo-label learning [9, 8]. For image translation based UDA [54], some researchers [55, 8] adopted Cycle-GAN [57] to generate translation images to reduce domain discrepancies. Zhang et al. [5] utilized Fourier transform to create pseudo-SAR images for alignment, while others [6] added speckle noise to optical images to mitigate domain discrepancies. Adversarial adaptation aligns domain features via adversarial learning [3]. Some efforts integrate adversarial adaptation with feature decoupling [11] or context relationships [10] to improve knowledge transfer. Liu et al. [56] leverage semantic calibration and adversarial feature alignment to enhance cross-domain ship detection. To address feature discriminability issues of adversarial learning, some efforts [8, 9] introduced pseudo-label learning strategies to mitigate feature degradation, while Dong et al. [58] leveraged attention mechanisms to improve transfer performance. Despite these advancements, UDA still has limitations. On one hand, unconditioned blind adaptation of UDA may lead to distortion of domain structures. On the other hand, The discriminative capability of most UDA methods are overly biased to the source domain. Directly applying such biased knowledge to the target domain can result in unreliable predictions, and aligning these unreliable predictions may even cause risky semantic propagation.
III Detection Model Enhanced by Evidential Learning
@@
III-A Evidential Modeling for Detection Model
The detection branch $\mathcal{F}_{D}$ is responsible for class prediction $y$ and $box$ regression. Since the target class prediction satisfies the characteristics of a Multinomial distribution, the class probability $p$ can be modeled as parameters of the Multinomial distribution [17, 19], which is expressed as: $y\sim P(y\mid p)=Mul(y\mid p)$.
The evidential learning branch $\mathcal{F}_{E}$, as shown in Fig. 2, is primarily designed to enhance the model’s reliability and quantify uncertainty. By introducing the Dirichlet distribution, it models the class probability to capture the confidence in the model’s predictions. The Dirichlet distribution $Dir(p\mid\alpha)$, as a conjugate prior of the Multinomial distribution, is defined as:
@@ III-A Evidential Modeling for Detection Model
=$=\begin{cases}\frac{1}{\beta(\alpha)}\prod_{c=1}^{C}p_{c}^{\alpha_{c}-1},&\text{if }p\in S_{C},\\ 0,&\text{otherwise}.\end{cases}$
@@
[…]
III-B Model Optimization Using Source Data
@@
$\mathrm{KL}\left({Q\left({p\mid\alpha}\right)\|P\left({p\mid\alpha}\right)}\right)\approx\mathrm{KL}\left({Dir\left({p\mid\dot{\alpha}}\right)\|Dir\left({p\mid[1,...,1]}\right)}\right).$ (6)
@@
For parameter optimization of the target task detection branch $\mathcal{F}_{D}$, the target detection loss $\mathcal{L}_{\textit{S}}$ can be directly used to optimize the model. For the evidential learning branch $\mathcal{F}_{E}$, its parameters can be optimized by minimizing ${\mathcal{L}_{\textit{EVI}}}$. Therefore, the total supervised loss for the source domain is presented as:
@@ IV Structural Hierarchy Adaptation and Reliable Adjacency Alignment
IV-A Scattering Distance Measurement
@@
Given a SAR target image $I_{i}$, the distribution of its strong scattering points is represented as: ${P_{i}}=\left\{{{x_{n}},{y_{n}},{a_{n}}}\right\}_{n=1}^{N}$. $(x_{n},y_{n})$ and $a_{n}$ respectively represent the spatial coordinates and intensities of scattering points in the image. Similarly, the scattering point set of another SAR target $I_{j}$ can be expressed as ${Q_{j}}=\left\{{{x_{m}},{y_{m}},{a_{m}}}\right\}_{m=1}^{M}$. Considering the sparsity and uneven distribution of scattering points, traditional distance measurement methods cannot effectively quantify such distribution differences. In contrast, based on optimal transport theory, Earth Mover’s Distance (EMD) [59] constructs the distribution difference between point sets $P_{i}$ and $Q_{j}$ as the minimum transportation cost from distribution $P_{i}$ to $Q_{j}$. EMD is robust to variations in the size, distribution, and domain of the scattering point sets. It can effectively measure the global and local differences between sparse scattering point sets under different resolutions and scales. The calculation method for scattering structure distance measurement is shown in Alg. 1.
@@
$f_{n,m}\geq 0,\quad\sum_{n=1}^{N}\sum_{m=1}^{M}f_{n,m}=1,$
$\sum_{m=1}^{M}f_{n,m}=w_{n},\quad\sum_{n=1}^{N}f_{n,m}=w_{m};$

### v2

@@ Abstract
In recent years, continuous improvements in SAR resolution have significantly benefited applications such as urban monitoring and target detection. However, these improvements in resolution have also led to increased discrepancies in scattering characteristics, posing challenges to the generalization ability of target detection models. While domain adaptation technologies provide a potential solution, the inevitable discrepancies caused by resolution differences often result in blind feature adaptation and unreliable semantic propagation, ultimately degrading the domain adaptation performance. To address these challenges, this paper proposes a novel SAR target detection method, termed CR-Net, which incorporates structure priors and evidential learning theory into the detection model, enabling reliable domain adaptation for cross-resolution detection. To be specific, CR-Net integrates Structure-induced Hierarchical Feature Adaptation (SHFA) and Reliable Structural Adjacency Alignment (RSAA). The SHFA module is designed to establish structural correlations between targets and achieve structure-aware feature adaptation, thereby enhancing the interpretability of the adaptation process. Afterwards, the RSAA module is proposed to enhance reliable semantic alignment, by leveraging the secure adjacency set to transfer valuable discriminative knowledge from the source domain to the target domain. This further improves the discriminability of the detection model in the target domain. Based on experimental results from different-resolution datasets, the proposed CR-Net significantly enhances cross-resolution adaptation by preserving intra-domain structures and improving discriminability. It achieves state-of-the-art (SOTA) performance in cross-resolution SAR target detection.
@@
Cross-resolution detection, evidential learning, reliable adjacency alignment, SAR target detection, scattering structure, unsupervised domain adaptation.
@@ I Introduction
Synthetic Aperture Radar (SAR) achieves high-resolution imaging by actively transmitting and receiving electromagnetic waves. This unique capability allows SAR to operate effectively under adverse weather conditions and regardless of daylight, enabling all-weather and all-day imaging. These attributes have made SAR widely applicable in the field of Earth observation [1, 2]. As one of the primary means of Earth observation, SAR target detection has been a hot topic in both research and applications. In recent years, advancements in SAR sensor technology have significantly improved the spatial resolution of SAR images, transitioning from traditional meter-level resolution to sub-meter resolution. This leap forward has unlocked significant potential for more refined Earth observation applications: sub-meter resolution SAR images can capture more target details, offering substantial advantages in fine-grained tasks such as urban monitoring [1], disaster assessment [3], and target recognition [4].
@@
Domain adaptation is an effective approach to address the aforementioned domain discrepancies and improve the generalization performance on the unseen data [5, 6]. The goal of this technique is to transfer the knowledge learned from the labeled source domain to an unlabeled target domain with distribution differences, by enhancing the model’s transferability and discriminability to gain satisfactory performance on the target domain data.
In recent years, some advanced researches have leveraged domain adaptation techniques in SAR target detection, focusing on improving cross-modality (e.g. optical-to-SAR [7, 8, 9] and SAR-to-SAR [10, 11, 12]) detection performance. For instance, Zhao et al. [13, 14] achieved SAR-to-SAR target detection by employing adversarial learning and transferable feature decomposition to extract domain-invariant features. However, such blind adaptation methods do not explicitly consider intra-domain variations, which may lead to feature misalignment and structural distortions within domains. To address this issue, Pan et al.[7] proposed an unbalanced adversarial feature alignment approach, enabling ship detection from optical to SAR images. Nevertheless, some studies[15, 16] have indicated that adversarial learning-based domain adaptation methods negatively affect feature discriminability. To mitigate this problem, several works [16, 11, 12] have incorporated pseudo-label learning to enhance feature discriminability. However, unreliable predictions in the target domain may lead to risky semantic propagation and error accumulation during the pseudo-label learning process. To reduce domain discrepancies between different sensors, other studies [10, 3, 17] have leveraged the physical priors of SAR targets to enhance transferability. Wang et al.[17] utilized sub-aperture decomposition[18] and data augmentation to achieve knowledge transfer from synthetic data to real-world data. Some researchers [10] have introduced key scattering points of targets to enable center-aware feature adaptation, while Zhang et al. [3] incorporated Fourier transform features [19] to establish local structural relationships and enable target detection from optical to SAR images. However, the feasibility of these methods for cross-resolution detection in SAR images has not yet been verified.
Although the aforementioned methods have made some progress in cross-domain SAR target detection, applying these black-box style models to cross-resolution SAR target detection still faces several critical challenges:
@@
    A Scattering Structure Distance (SSD) is proposed to quantify the structure similarity between SAR targets. Leveraging the scattering structure similarity of instances, a Structure-induced Hierarchical Feature Adaptation (SHFA) module is developed. By introducing structural constraints, SHFA module facilitates hierarchical feature adaptation between the source and target instances, enhancing the feature transferability and avoiding structural distortions throughout the adaptation process.
@@
    A Reliable Structural Adjacency Alignment (RSAA) module is proposed to capture reliable adjacency relationships by aligning semantic information between reliable neighbors across domains. RSAA module effectively transfers discriminative semantics from the source domain to the target domain, significantly improving the model’s performance on the target domain while mitigating error accumulation.
@@
    The proposed CR-Net combines SHFA and RSAA modules, achieving significant performance improvements in cross-resolution SAR target detection tasks. Experimental results demonstrate that CR-Net not only preserves the intra-domain structures but also significantly enhances the reliability and accuracy of cross-resolution detection.
The remainder of this paper is organized as follows. Section II reviews related works. Section III explains how evidence learning is incorporated into SAR detection models. Building on the evidence-enhanced SAR detection framework, Section IV details the proposed cross-resolution detection method. Section V presents the experimental results and analysis. Finally, Section VI concludes the paper.
@@ II-A Uncertainty Estimation and Evidential Learning
Uncertainty is a key factor for evaluating the predictive reliability of deep models [20, 21]. Bayesian deep learning [22] and subsequent approximate methods [23, 24] provide theoretical and practical approaches for uncertainty modeling. However, they still face challenges related to high computational cost and low efficiency. Recently, evidence learning [20] has gained significant attention due to its capability to explicitly model and inference with evidence, making it a powerful approach for quantifying uncertainty. Compared to traditional Bayesian methods, evidence learning learns a Dirichlet distribution to simultaneously output predictive confidence and uncertainty, avoiding high computational costs and improving flexibility [25]. Currently, evidential learning has been widely applied to high-reliability tasks such as medical diagnosis [26] and anomaly detection [27, 28]. However, its application in cross-resolution SAR target detection remains largely unexplored. This paper applies evidence learning in cross-resolution SAR detection tasks to model uncertainty and improve adaptation reliability across resolutions.
@@ II-B SAR Target Detection
In SAR target detection, mainstream approaches are typically divided into one-stage detectors (e.g., RetinaNet [29], YOLOs [30]) and two-stage detectors (e.g., Faster RCNN [31], Sparse RCNN [32]). Building on these frameworks, the SAR community has proposed various advancements. For example, Zhou et al. [33] developed a sidelobe-aware network for better small ship detection, while Ju et al. [34] and Zhao et al. [35] improved rotated target detection through polar encoding and enhanced inter-class separability [36]. Yang et al. [37] proposed a dynamic detection framework to boost dense prediction performance in complex scenarios. For two-stage detectors, Cui et al. [38] incorporated attention mechanisms [39] for improved multi-scale detection, and Cheng et al. [40] addressed nearshore detection with a saliency enhancement algorithm. Some studies [41, 42] have leveraged generative adversarial networks [43] (GANs) to enhance model discrimination and robustness. For instance, Ren et al. [41] introduced a Confucius tri-learning framework, which jointly trains two classifiers and a generator to achieve better discrimination performance with limited samples. Ju et al. [42] employed GANs to suppress background interference, thus improving detection in complex environments. Some researchers [44] utilized GANs to guide models in extracting essential structural features, thereby boosting discriminative capability. Other works [45, 46, 18] have focused on improving detection performance by mitigating environmental clutter [45], multi-scale dynamic learning [46], and optimizing data quality with data transformations or sub-aperture decomposition [47, 18].
Although these detection methods perform well when the training and testing data are independently and identically distributed, their performance drops significantly when confronted with data from different sensors, modalities, or resolutions.
@@ II-C Unsupervised Domain Adaptation
Unsupervised Domain Adaptation (UDA) is an important research direction in transfer learning, aiming to improve the performance of models on the target domain by leveraging labeled data from the source domain and unlabeled data from the target domain [5]. The core challenge lies in the distribution discrepancy (i.e., domain shift) between the source and target domains. Researchers have proposed various methods to reduce this discrepancy, including distribution alignment (e.g., DAN [5] and MADA [48]), pseudo-label generation [49, 16], image translation methods [15, 50], and mean-teacher training methods [51, 52]. UDA is widely applied in fields such as computer vision [53, 54, 55], medical image analysis [56], and autonomous driving [52]. Despite significant progress, UDA still faces challenges such as interpretability [57, 58], domain structure distortion [5], and unreliable knowledge transfer [21].
@@ II-D UDA for SAR Target Detection
In recent years, UDA has become a popular research topic in SAR target detection. Mainstream methods include image translation [15, 59], adversarial adaptation [60, 61], and pseudo-label learning [12, 11]. Image translation-based UDA [59] is a two-stage approach. Researchers have adopted CycleGAN [62, 60, 11] to generate translated images that reduce domain discrepancies, followed by domain adaptation techniques to enhance detection performance. Zhang et al. [8] created pseudo-SAR images using the Fourier transform for feature alignment, while others [9] added speckle noise to optical images for further mitigate domain discrepancies. To improve detection on low-quality SAR images, Pu et al. [63] employed CycleGAN for image enhancement, thereby boosting target detection accuracy. However, as these two-stage approaches are typically not trained in an end-to-end manner, it is difficult to guarantee optimal performance. In contrast, adversarial adaptation aligns domain feature distributions via end-to-end adversarial learning [6]. Some efforts integrate adversarial adaptation with feature decoupling [14] or context relationships [13] to facilitate knowledge transfer. Liu et al. [61] leverage semantic calibration and adversarial feature alignment to enhance cross-domain ship detection. However, some studies have revealed that such blind adversarial training may lead to distortion of domain structures and impair the discriminative capability of learned features. To alleviate the degradation of feature discriminability caused by adversarial learning, some works [12, 11, 60] have proposed incorporating pseudo-label learning strategies into the domain adaptation process. However, these methods essentially rely on the discriminative capability learned from the source domain data to select pseudo-labels of the target domain data. As a result, the discriminative capability of such UDA methods remains overly biased toward the source domain. Directly applying such biased knowledge to the target domain can result in unreliable predictions, and aligning these unreliable predictions may even cause risky semantic propagation and error accumulation.
III Preliminaries
@@
III-A Detection Model Enhanced by Evidential Learning
The detection branch $\mathcal{F}_{D}$ is responsible for class prediction $y$ and boxbox regression. Since the target class prediction satisfies the characteristics of a Multinomial distribution, the class probability $p$ can be modeled as parameters of the Multinomial distribution [20, 21], which is expressed as: $y\sim P(y\mid p)=Mul(y\mid p)$.
The evidential learning branch $\mathcal{F}_{E}$, as shown in Fig. 2, is primarily designed to enhance the model’s reliability and quantify uncertainty. By introducing the Dirichlet distribution, the model can represent class probabilities in a way that captures the confidence in its predictions. The Dirichlet distribution $Dir(p\mid\alpha)$, as a conjugate prior of the Multinomial distribution, is defined as:
@@ III-A Detection Model Enhanced by Evidential Learning
[…]
@@
where $S=\sum\nolimits_{c=1}^{C}{{e_{c}}}+C$, representing the Dirichlet strength. When the belief mass $b_{c}$ is larger, it indicates that more evidence $e_{c}$ is assigned to class $c$. The discriminative probability of $c\text{-th}$ class is $p_{c}=\alpha_{c}/S$. A larger Dirichlet strength $S$ reduces the overall uncertainty $u$, increasing the confidence and reliability in class predictions.
III-B Model Optimization Using Source Domain Data
@@
$\mathrm{KL}(Q(p\mid\alpha)$ $\|\ P(p\mid\alpha))$ (6)
$\approx\ \mathrm{KL}\left({Dir}(p\mid\dot{\alpha})\ \|\ {Dir}(p\mid[1,\dots,1])\right)$
@@
For parameter optimization of the target task detection branch $\mathcal{F}_{D}$, the target detection loss $\mathcal{L}_{\textit{S}}$ is directly used to optimize the model. For the evidential learning branch $\mathcal{F}_{E}$, its parameters are optimized by minimizing $\mathcal{L}_{\textit{EVI}}$}. Therefore, the total supervised loss for the source domain is presented as:
@@ IV Structural Hierarchy Adaptation and Reliable Adjacency Alignment
IV-A Scattering Structure Distance
@@
Given a SAR target image $I_{i}$, the distribution of its strong scattering points is represented as: ${P_{i}}=\left\{{{x_{n}},{y_{n}},{a_{n}}}\right\}_{n=1}^{N}$. $(x_{n},y_{n})$ and $a_{n}$ respectively represent the spatial coordinates and intensities of scattering points in the image. Similarly, the scattering point set of another SAR target $I_{j}$ can be expressed as ${Q_{j}}=\left\{{{x_{m}},{y_{m}},{a_{m}}}\right\}_{m=1}^{M}$.
Considering the sparsity and uneven distribution of scattering points, traditional distance measurement methods cannot effectively quantify such distribution differences. In contrast, Earth Mover’s Distance (EMD) [64], based on optimal transport theory [65, 66], provides a more effective way to measure the distribution discrepancy between scattering point sets. Specifically, EMD defines the distance between point sets $P_{i}$ and $Q_{j}$ as the minimum transportation cost required to transform the distribution $P_{i}$ into $Q_{j}$. It is robust to variations in the size and distribution, and it also can effectively capture both global and local differences between sparse scattering point sets under different resolutions and scales. Therefore, we propose to use EMD for calculating the scattering structure distance (SSD) between point sets.
The detailed calculation method for SSD is presented in Alg. 1. Specifically, SSD is computed by solving for the minimum transportation cost based on EMD. In the SSD formulation, $f_{n,m}$ denotes the flow transported from the $n$-th point in $P_{i}$ to the $m$-th point in $Q_{j}$, while $w_{n}$ and $w_{m}$ represent the normalized weights (masses) assigned to the $n$-th source point and $m$-th target point, respectively. By optimizing the flows $f_{n,m}$ with respect to the sum constraints defined by $w_{n}$ and $w_{m}$, the SSD is computed as the resulting minimum transportation cost. To evaluate its effectiveness, Fig. 5 compares the discriminability of different distance metrics across various target structures. Compared with traditional metrics such as L2 distance, cosine distance, and Hausdorff distance, the proposed SSD achieves a significantly lower distance distribution for intra-class target structure pairs (e.g., A220 vs. A220) and a higher separation from inter-class pairs (e.g., A220 vs. A330), resulting in less overlap between categories and greater inter-group discriminability. In addition, SSD yields a more compact and less dispersed intra-class distance distribution, indicating improved stability. These results highlight the effectiveness of SSD in handling discrete and uneven scattering points, and also demonstrate its clear advantages in distinguishing between different SAR target structures.
@@
$f_{n,m}\geq 0,\quad\sum_{n=1}^{N}\sum_{m=1}^{M}f_{n,m}=1;$
$\sum_{m=1}^{M}f_{n,m}=w_{n},\quad\forall n=1,\ldots,N;$}
$\sum_{n=1}^{N}f_{n,m}=w_{m},\quad\forall m=1,\ldots,M;$}
@@
Figure 5: Boxplots of different distance metrics for SAR target structures. (a)-(d) correspond to L2 distance, cosine distance, Hausdorff distance, and the proposed scattering structure distance (SSD), respectively. Compared with other distance measures, SSD shows superior discriminative ability for structural differences among different target categories.